# Leptoquark-induced CLFV decays with a light SM-singlet scalar

### v2

@@ Abstract
The Standard Model (SM), if augmented with a light SM-singlet scalar $\phi$, 2-body charged lepton flavor violating (CLFV) decay channels can be accessed with $\phi$ as one of the final states. The effective ineteractions between $\phi$ and the SM fields can arise at TeV-scale where the governing theory extends the SM particle spectrum with a scalar leptoquark (LQ) $S_{1}$. Further, in the presence of $S_{1}$, $\phi$ can mediate 3-body CLFV processes with either two photons or two gluons in the final states. Thus, the model predicts an exotic 3-body CLFV channel: $\ell_{A}\to\ell_{B}gg$, which can be tested/constrained only through some future high-energy experiments looking for di-gluon signal from a leptonic decay.
@@ I Introduction
Experimental searches for CLFV processes have been going on for a long Hincks:1948vr ; PhysRev.74.1364 , with the current sensitivity reaching $\sim\mathcal{O}(10^{-13})$ for a few flavor violating channels MEG:2016leq ; SINDRUM:1987nra ; SINDRUMII:2006dvw and being significantly improved for the others. However, in the absence of any positive signal, interest has grown in searching for CLFV decays involving low-energy BSM particles. For example, Belle-II presents the best current upper limit on $\tau\to e\phi$ and $\tau\to\mu\phi$ at 95% C.L. Belle-II:2022heu , where $\phi$ is an invisible spin-0 boson in the mass range $0-1.6$ GeV. Similar bounds for $\mu\to e\phi$ is given by Refs. TWIST:2014ymv ; Jodidio:1986mz . However, if $\phi$ decays to visible particles inside the detector, more stringent constraints could be placed from the non-observation of 3-body CLFV processes.
@@
Various models have been proposed to address the current bounds on $\ell_{A}\to\ell_{B}\phi$, with $\phi$ being either a light scalar/pseudoscalar Wilczek:1982rv ; Grinstein:1985rt ; Berezhiani:1989fp ; Feng:1997tn ; Hirsch:2009ee ; Jaeckel:2013uva ; Celis:2014iua ; Celis:2014jua ; Galon:2016bka ; Calibbi:2016hwq ; Ema:2016ops ; Bjorkeroth:2018dzu ; Bauer:2019gfk ; Heeck:2019guh ; Cornella:2019uxs ; Calibbi:2020jvd ; Escribano:2020wua or a gauge boson Farzan:2015hkd ; Heeck:2016xkh ; Farzan:2016wym ; Ibarra:2021xyk , associated with some spontaneously broken symmetry. This paper considers $\phi$ to be a generic real SM-singlet scalar where the flavor-violating interactions arise at one-loop level in the presence of a scalar LQ $S_{1}$. LQs (for a review, see Ref. Dorsner:2016wpm ) appear naturally in the grand unified theories (GUT) Pati:1974yy ; Georgi:1974sy ; Georgi:1974my and work as an excellent BSM candidate to induce flavor violation. Though in the simplest GUT models LQs acquire mass at a scale $\Lambda_{\rm GUT}\gg\mathcal{O}(1)$ TeV, there exist GUT formulations which can ensure the stability of proton with a TeV-scale scalar LQ BUCHMULLER1986377 ; Murayama:1991ah ; Dorsner:2005fq ; GEORGI1979297 ; FileviezPerez:2007bcw ; Senjanovic:1982ex . Thus, the present paper, while extending the SM with a TeV-scale scalar LQ to describe the interactions between $\phi$ and the SM fields, effectively portrays a UV-complete theory. Ref. Mandal:2019gff has already elaborated on various CLFV observables arising in the presence of a TeV-scale scalar LQ and used the experimental upper limits to constrain the New Physics (NP) couplings. However, interesting consequences may appear if one considers a light SM-singlet scalar within the $S_{1}$-extension of SM. For example, with $M_{\phi}<(m_{\mu}-m_{e})$, $\phi$ can be produced through the on-shell decays of $\tau$ and $\mu$ leading to the CLFV processes $\tau\to\ell_{B}\phi$ [$\ell_{B}=e,\,\mu$] and $\mu\to e\phi$, respectively. In practice, the term light will be used in this paper to refer to $M_{\phi}<2m_{e}$, i.e., a sub-MeV gauge-singlet scalar. In this mass regime, it can be easily ensured that the di-lepton and di-quark decay modes of $\phi$ De:2024tbo are kinematically forbidden. However, $\phi$ can always decay to two photons at one-loop level, leading to 3-body CLFV channels of the form $\ell_{A}\to\ell_{B}\gamma\gamma$. Thus, the experimental bounds for $\ell_{A}\to\ell_{B}\gamma\gamma$ can be used to check the consistency of the parameter space allowed through the $\ell_{A}\to\ell_{B}\phi$ searches. Further, LQs being colored particles, within this proposed BSM formulation $\phi$ can also decay to two gluons. This results in a rarely discussed unconstrained CLFV decay process: $\ell_{A}\to\ell_{B}gg$. Indeed, this is a significant phenomenological outcome of this model as it predicts new CLFV observables. Therefore, any future search aiming for $\ell_{A}\to\ell_{B}gg$ will be crucial to test/falsify the model.
@@ III Lepton Flavor Violation: $\boldsymbol{\Large\ell_{A}\to\ell_{B}+\phi}$
Eq. (5) has been obtained assuming an on-shell decay of $\ell_{A}$ so that $p_{A}^{2}=(p_{B}+p_{\phi})^{2}=m_{A}^{2}$. Here, $C_{0}$ and $D_{0}$ denote the standard Passarino-Veltmen functions Romao:2020 ; Hahn:1998yk for the scalar 3-point and 4-point one-loop integrals, respectively.
@@ IV Numerical Analysis
As stated in Sec. III, the $S_{1}$-extension of SM can accommodate several BSM observations and possibilities resulting in strong bounds on the NP couplings $\xi^{ij}_{L,R}$. Thus, to analyze the $\ell_{A}\to\ell_{B}\phi$ processes, one must choose a parameter space consistent with all the existing leptonic constraints. Following Ref. Mandal:2019gff , Table 2 enlists the most stringent upper limits on $|\Omega^{AB}_{i}|^{2}$ relavent for the present study.
@@
The first column of Table 2 specifies a particular $\ell_{A}\to\ell_{B}\phi$ process, while the third column represents the experimental bounds on the associated NP couplings arising through the other CLFV processes. However, the bounds drastically deteriorate with increasing LQ mass. Thus, in general, the couplings considered for the present analysis assume $M_{S_{1}}\geq 2$ TeV and have been listed in the last column of Table 2. As an illustrative choice, $\beta=5$ TeV will be used for the analysis.
@@ V.1 $\boldsymbol{\ell_{A}\to\ell_{B}\gamma\gamma}$
Figure 7: Variation of (a) BR$(\tau\to e\gamma\gamma)$, (b) BR$(\tau\to\mu\gamma\gamma)$, and (c) BR$(\mu\to e\gamma\gamma)$ as a function of $M_{\phi}$ for $M_{S_{1}}=2$ TeV (red), 3 TeV (sky), and 4 TeV (yellow). The black solid lines represent the corresponding experimental sentitivities (at 90% C.L.).
This type of 3-body CLFV processes have been searched for a long time (particularly for $\mu\to e\gamma\gamma$) through various experiments, and till now, no flavor violation has been observed. Thus, the null results have set certain upper limits on BR($\ell_{A}\to\ell_{B}\gamma\gamma$), which may further propagate into the $(\ell_{A}\to\ell_{B}\phi)$-allowed parameter space as more stringent constraints. Table 3 presents the current experimental bounds on BR($\ell_{A}\to\ell_{B}\gamma\gamma$).
@@ V.2 $\boldsymbol{\ell_{A}\to\ell_{B}gg}$
The proposed extension of SM can accommodate an exotic 3-body CLFV channel where the heavier charged leptons can decay to the lighter ones and two gluons. To the best of the author’s knowledge, very little has been explored about this particular type of CLFV decays, as they can never be probed through low-energy experiments. Colliders using deep inelastic scattering can only be an option to search for $\ell_{A}\to\ell_{B}gg$ AbdulKhalek:2022hcn . For example, Refs. Takeuchi:2017btl ; Cirigliano:2021img have discussed the gluonic contribution to the $\tau$-specific CLFV processes in an Electron Ion/Proton Collider.
Therefore, these decays might lead to completely new observables for probing the lepton flavor violation in the charged sector. Fig. 8 shows the model predictions for $\tau\to egg$, $\tau\to\mu gg$, and $\mu\to egg$. Note that, ${\rm BR}(\ell_{A}\to\ell_{B}gg)/{\rm BR}(\ell_{A}\to\ell_{B}\gamma\gamma)\sim\mathcal{O}(10^{3})$ as $\phi$ dominantly decays through the di-gluon channel. Therefore, in principle, $\ell_{A}\to\ell_{B}gg$ should have a better detection prospect compared to $\ell_{A}\to\ell_{B}\gamma\gamma$ if the background is substantially reduced ATL-PHYS-PUB-2017-017 ; CMS-DP-2017-027 ; Andrews:2019faz . Though, currently, there is no direct experimental bound for these processes, detectors looking for di-gluon signals from CLFV decays may test/falsify the present predictions in the future.
@@ VI Conclusion
By the next decade, experiments searching for CLFV processes is going to play a vital role in discovering physics beyond the SM. The chances will definitely enhance with the increasing number CLFV observables. Moreover, the observation will also be significant in understanding the possible NP theory associated with such a positive signal. This paper has extended the SM with a light SM-singlet scalar $\phi$ with mass $M_{\phi}<2m_{e}$. There is a wide range of theories where such a generic scalar can be associated with a spontaneously broken global symmetry. The effective couplings of $\phi$ with the SM fields can be obtained at one-loop level in the presence of a TeV-scale scalar LQ $S_{1}$. Phenomenologically, scalar LQs are well-motivated hypothetical particles that generate a large variety of CLFV processes. Thus, the non-observation of any CLFV process with the existing experimental sensitivities results in stringent constraints on the NP couplings at the quark-LQ-lepton vertices. Considering such a pre-constrained parameter space, the present paper has studied the discovery prospects of $\tau\to e\phi$, $\tau\to\mu\phi$, and $\mu\to e\phi$. Though for illustrations, 1 keV $\leq M_{\phi}\leq$ 1 MeV has been chosen, the results are equally valid for any generic scalar lighter than $2m_{e}$. $S_{1}$ being an EM-charged color-triplet scalar, $\phi$ can decay through di-photon and di-gluon channels. This induces the possibility of 3-body CLFV decays $\ell_{A}\to\ell_{B}\gamma\gamma$ and $\ell_{A}\to\ell_{B}gg$, where $\phi$ acts as the mediator. Though for the $\tau$-sector, the experimental bounds on $\tau\to\ell_{B}\gamma\gamma$ [$\ell_{B}=e,\,\mu$] are less stringent compared to those on the $\tau\to\ell_{B}\phi$, for $\mu$-sector BR$(\mu\to e\gamma\gamma)$ becomes crucial to constrain the parameter space. The most remarkable feature of this model is the possibility to predict CLFV observables corresponding to the 3-body decay $\ell_{A}\to\ell_{B}gg$, for all the possible channels. No direct experimental limit exists for such leptonic decays with two gluons in the final states. Using narrow-width approximation, the model has predicted for BR$(\tau\to egg)$, BR$(\tau\to\mu gg)$, and BR$(\mu\to egg)$ within the allowed parameter space. Therefore, a future experiment searching for $\ell_{A}\to\ell_{B}gg$ will be significant to test/constrain the model or may lead to a discovery.

### v3

@@ Abstract
The Standard Model (SM), if augmented with a light SM-singlet scalar $\phi$ and a TeV-scale scalar leptoquark (LQ) $S_{1}$, 2-body charged lepton flavor violating (CLFV) decay channels can be accessed with $\phi$ as one of the final states, where the leading order effective interactions between $\phi$ and the SM fields arise at one-loop level. Further, in the presence of $S_{1}$, $\phi$ can mediate 3-body CLFV processes with either two photons or two gluons in the final states. Thus, the model predicts an exotic 3-body CLFV channel: $\ell_{A}\to\ell_{B}gg$, which can be tested/constrained only through some future high-energy experiments looking for di-gluon signals from a leptonic decay.
@@ I Introduction
Experimental searches for CLFV processes have been going on for a long time Hincks:1948vr ; PhysRev.74.1364 , with the current sensitivity reaching $\sim\mathcal{O}(10^{-13})$ for a few flavor violating channels MEG:2016leq ; SINDRUM:1987nra ; SINDRUMII:2006dvw and being significantly improved for the others. However, in the absence of any positive signal, interest has grown in searching for CLFV decays involving low-energy BSM particles. For example, Belle-II presents the best current upper limit on $\tau\to e\phi$ and $\tau\to\mu\phi$ at 95% C.L. Belle-II:2022heu , where $\phi$ is an invisible spin-0 boson in the mass range $0-1.6$ GeV. Similar bounds for $\mu\to e\phi$ are given by Refs. TWIST:2014ymv ; Jodidio:1986mz . However, if $\phi$ decays to visible particles inside the detector, more stringent constraints could be placed from the non-observation of 3-body CLFV processes.
@@
Various models have been proposed to address the current bounds on $\ell_{A}\to\ell_{B}\phi$, with $\phi$ being either a light scalar/pseudoscalar PhysRevLett.48.11 ; Wilczek:1982rv ; Grinstein:1985rt ; Berezhiani:1989fp ; Feng:1997tn ; Hirsch:2009ee ; Jaeckel:2013uva ; Celis:2014iua ; Celis:2014jua ; Galon:2016bka ; Calibbi:2016hwq ; Ema:2016ops ; Bjorkeroth:2018dzu ; Bauer:2019gfk ; Heeck:2019guh ; Cornella:2019uxs ; Calibbi:2020jvd ; Escribano:2020wua or a gauge boson Farzan:2015hkd ; Heeck:2016xkh ; Farzan:2016wym ; Ibarra:2021xyk , associated with some spontaneously broken symmetry. This paper considers $\phi$ to be a generic real SM-singlet scalar where the flavor-violating interactions arise at one-loop level in the presence of a scalar LQ $S_{1}$. LQs (for a review, see Ref. Dorsner:2016wpm ) appear naturally in the grand unified theories (GUT) Pati:1974yy ; Georgi:1974sy ; Georgi:1974my and work as an excellent BSM candidate to induce flavor violation. Though in the simplest GUT models LQs acquire mass at a scale $\Lambda_{\rm GUT}\gg\mathcal{O}(1)$ TeV, there exist GUT formulations that can ensure the stability of proton with a TeV-scale scalar LQ Dorsner:2004jj ; Aydemir:2019ynb ; Dorsner:2024seb . Thus, the present paper, while extending the SM with a TeV-scale scalar LQ to describe the interactions between $\phi$ and the SM fields, effectively portrays a UV-complete theory. Ref. Mandal:2019gff has already elaborated on various CLFV observables arising in the presence of a TeV-scale scalar LQ and used the experimental upper limits to constrain the New Physics (NP) couplings. However, interesting consequences may appear if one considers a light SM-singlet scalar within the $S_{1}$-extension of SM. For example, with $M_{\phi}<(m_{\mu}-m_{e})$, $\phi$ can be produced through the on-shell decays of $\tau$ and $\mu$ leading to the CLFV processes $\tau\to\ell_{B}\phi$ [$\ell_{B}=e,\,\mu$] and $\mu\to e\phi$, respectively. In practice, the term light will be used in this paper to refer to $M_{\phi}<2m_{e}$, i.e., a sub-MeV gauge-singlet scalar. In this mass regime, it can be easily ensured that the di-lepton and di-quark decay modes of $\phi$ De:2024tbo are kinematically forbidden. However, $\phi$ can always decay to two photons at one-loop level, leading to 3-body CLFV channels of the form $\ell_{A}\to\ell_{B}\gamma\gamma$. Thus, the experimental bounds for $\ell_{A}\to\ell_{B}\gamma\gamma$ can be used to check the consistency of the parameter space allowed through the $\ell_{A}\to\ell_{B}\phi$ searches. Further, LQs being colored particles, within this proposed BSM formulation $\phi$ can also decay to two gluons. This results in a rarely discussed unconstrained CLFV decay process: $\ell_{A}\to\ell_{B}gg$. Indeed, this is a significant phenomenological outcome of this model as it predicts a set of unexplored CLFV observables. Therefore, any future search aiming for $\ell_{A}\to\ell_{B}gg$ will be crucial to test/falsify the model.
@@ III Lepton Flavor Violation: $\boldsymbol{\Large\ell_{A}\to\ell_{B}+\phi}$
Eq. (5) has been obtained assuming an on-shell decay of $\ell_{A}$ so that $p_{A}^{2}=(p_{B}+p_{\phi})^{2}=m_{A}^{2}$. Here, $C_{0}$ and $D_{0}$ denote the standard Passarino-Veltman functions Romao:2020 ; Hahn:1998yk for the scalar 3-point and 4-point one-loop integrals, respectively.
@@ IV Numerical Analysis
As stated in Sec. III, the $S_{1}$-extension of SM can accommodate several BSM observations and possibilities resulting in strong bounds on the NP couplings $\xi^{ij}_{L,R}$. Thus, to analyze the $\ell_{A}\to\ell_{B}\phi$ processes, one must choose a parameter space consistent with all the existing leptonic constraints. Following Ref. Mandal:2019gff , Table 2 enlists the most stringent upper limits on $|\Omega^{AB}_{i}|^{2}$ relevant for the present study.
@@
The first column of Table 2 specifies a particular $\ell_{A}\to\ell_{B}\phi$ process, while the third column represents the experimental bounds on the associated NP couplings arising through the other CLFV processes. However, the bounds drastically deteriorate with increasing LQ mass. In general, the couplings considered for the present analysis assume $M_{S_{1}}\geq 2$ TeV ParticleDataGroup:2022pth and have been listed in the last column of Table 2. As an illustrative choice, $\beta=5$ TeV will be used for the analysis.
@@ V.1 $\boldsymbol{\ell_{A}\to\ell_{B}\gamma\gamma}$
Figure 7: Variation of (a) BR$(\tau\to e\gamma\gamma)$, (b) BR$(\tau\to\mu\gamma\gamma)$, and (c) BR$(\mu\to e\gamma\gamma)$ as a function of $M_{\phi}$ for $M_{S_{1}}=2$ TeV (red), 3 TeV (sky), and 4 TeV (yellow). The black solid lines represent the corresponding experimental sensitivities (at 90% C.L.).
This type of 3-body CLFV processes has been searched for a long time (particularly for $\mu\to e\gamma\gamma$) through various experiments, and till now, no flavor violation has been observed. Thus, the null results have set certain upper limits on BR($\ell_{A}\to\ell_{B}\gamma\gamma$), which may further propagate into the $(\ell_{A}\to\ell_{B}\phi)$-allowed parameter space as more stringent constraints. Table 3 presents the current experimental bounds on BR($\ell_{A}\to\ell_{B}\gamma\gamma$).
@@ V.2 $\boldsymbol{\ell_{A}\to\ell_{B}gg}$
The proposed extension of SM can accommodate an exotic 3-body CLFV channel where the heavier charged leptons can decay to the lighter ones and two gluons. To the best of the author’s knowledge, very little has been explored about this particular type of CLFV decays, as they can never be probed through low-energy experiments. Colliders using deep inelastic scattering can only be an option to search for $\ell_{A}\to\ell_{B}gg$ AbdulKhalek:2022hcn . For example, Refs. Takeuchi:2017btl ; Cirigliano:2021img have discussed the gluonic contribution to the $\tau$-specific CLFV processes in an Electron-Ion/Proton Collider.
Therefore, these decays might lead to completely new observables for probing the lepton flavor violation in the charged sector. Fig. 8 shows the model predictions for $\tau\to egg$, $\tau\to\mu gg$, and $\mu\to egg$. Note that, ${\rm BR}(\ell_{A}\to\ell_{B}gg)/{\rm BR}(\ell_{A}\to\ell_{B}\gamma\gamma)\sim\mathcal{O}(10^{3})$ as $\phi$ dominantly decays through the di-gluon channel. Therefore, in principle, $\ell_{A}\to\ell_{B}gg$ should have a better detection prospect compared to $\ell_{A}\to\ell_{B}\gamma\gamma$ if the background is substantially reduced ATL-PHYS-PUB-2017-017 ; CMS-DP-2017-027 ; Andrews:2019faz . Though, currently, there is no direct experimental bound on these processes, detectors looking for di-gluon signals from CLFV decays may test/falsify the present predictions in the future.
@@ VI Conclusion
By the next decade, experiments searching for CLFV processes are going to play a vital role in discovering physics beyond the SM. The chances will definitely be enhanced with the increasing number of CLFV observables. Moreover, the observation will also be significant in understanding the possible NP theory associated with such a positive signal. This paper has extended the SM with a light SM-singlet scalar $\phi$ with mass $M_{\phi}<2m_{e}$ and a TeV-scale scalar LQ $S_{1}$. There is a wide range of theories where such a generic gauge-singlet scalar can be associated with a spontaneously broken global symmetry. The effective couplings of $\phi$ with the SM fields can be obtained at one-loop level in the presence of $S_{1}$. Phenomenologically, scalar LQs are well-motivated hypothetical particles that generate a large variety of CLFV processes. Thus, the non-observation of any CLFV process with the existing experimental sensitivities results in stringent constraints on the NP couplings at the quark-LQ-lepton vertices. Considering such a pre-constrained parameter space, the present paper has studied the discovery prospects of $\tau\to e\phi$, $\tau\to\mu\phi$, and $\mu\to e\phi$. Though for illustrations, 1 keV $\leq M_{\phi}\leq$ 1 MeV has been chosen, the results are equally valid for any generic scalar lighter than $2m_{e}$. $S_{1}$ being an EM-charged color-triplet scalar, $\phi$ can decay through di-photon and di-gluon channels. This induces the possibility of 3-body CLFV decays $\ell_{A}\to\ell_{B}\gamma\gamma$ and $\ell_{A}\to\ell_{B}gg$, where $\phi$ acts as the mediator. Though for the $\tau$-sector, the experimental bounds on $\tau\to\ell_{B}\gamma\gamma$ [$\ell_{B}=e,\,\mu$] are less stringent compared to those on the $\tau\to\ell_{B}\phi$, for $\mu$-sector BR$(\mu\to e\gamma\gamma)$ becomes crucial to constrain the parameter space. The most remarkable feature of this model is the possibility to predict CLFV observables corresponding to the 3-body decay $\ell_{A}\to\ell_{B}gg$, for all the possible channels. No direct experimental limit exists for such leptonic decays with two gluons in the final states. Using narrow-width approximation, the model has predicted for BR$(\tau\to egg)$, BR$(\tau\to\mu gg)$, and BR$(\mu\to egg)$ within the allowed parameter space. Therefore, a future experiment searching for $\ell_{A}\to\ell_{B}gg$ will be significant to test/constrain the model or may lead to a discovery.